# RBLA: Rank-Based-LoRA-Aggregation for Fine-tuning Heterogeneous Models in FLaaS

### v1

@@ Abstract
Federated Learning (FL) is a promising privacy-aware distributed learning framework that can be deployed on various devices, such as mobile phones, desktops, and devices equipped with CPUs or GPUs. In the context of server-based Federated Learning as a Service (FLaas), FL enables the central server to coordinate the training process across multiple devices without direct access to the local data, thereby enhancing privacy and data security. Low-Rank Adaptation (LoRA) is a method that fine-tunes models efficiently by focusing on a low-dimensional subspace of the model’s parameters. This approach significantly reduces computational and memory costs compared to fine-tuning all parameters from scratch. When integrated with FL, especially in a FLaas environment, LoRA allows for flexible and efficient deployment across diverse hardware with varying computational capabilities by adjusting the local model’s rank. However, in LoRA-enabled FL, different clients may train models with varying ranks, which poses a challenge for model aggregation on the server. Current methods of aggregating models of different ranks require padding weights to a uniform shape, which can degrade the global model’s performance. To address this issue, we propose Rank-Based LoRA Aggregation (RBLA), a novel model aggregation method designed for heterogeneous LoRA structures. RBLA preserves key features across models with different ranks. This paper analyzes the issues with current padding methods that reshape models for aggregation in a FLaas environment. Then, we introduce RBLA, a rank-based aggregation method that maintains both low-rank and high-rank features. Finally, we demonstrate the effectiveness of RBLA through comparative experiments with state-of-the-art methods.
@@ 1 Introduction
Neural Networks (NN) have become a widely applied approach in contemporary Computer Vision (CV) and Natural Language Processing (NLP). Traditionally, model training has involved collecting task-specific data and conducting centralized training in high-performance data centers. However, this centralized method raises significant privacy concerns, especially in applications dealing with sensitive data such as health information. Collecting and uploading user data to centralized servers inherently may lead to privacy breaches. To address these privacy challenges, McMahan et al. proposed Federated Learning (FL) in 2017 [22]. FL is a privacy-preserving distributed machine learning framework designed for decentralized neural network training. In this algorithm, models are trained locally on client devices (without sharing raw data), and model updates are then uploaded to a central server for aggregation. This concept has further evolved into Federated Learning as a Service (FLaaS) [15], where federated learning capabilities are provided as a cloud-based service. FLaaS simplifies the deployment process by managing the underlying infrastructure, allowing the implementation of federated learning across a diverse range of devices, including smartphones, IoT devices, and edge servers. This flexibility enables organizations to leverage privacy-preserving machine learning on a large scale without the complexities of building and maintaining their own FL systems. Despite these advantages, decentralized local model training in FL typically requires robust client device performance and communication capabilities. In practical FlaaS scenarios, the significant performance variations among client devices cause a challenge in deploying models with identical structures across all clients.
@@
Recent advancements in computing power and algorithms have led to numerous applications that use large models to process vast amounts of data on mobile devices, providing responses or making decisions. Examples include ChatGPT and Tesla’s autonomous driving technology. However, the inference and training processes of large models are complex and resource-intensive. Additionally, the heterogeneous nature of client devices in FL scenarios, with significant differences in performance, complicates the deployment and fine-tuning of federated large models. To address these issues, Low-Rank Adaptation (LoRA) [9] has emerged as a feasible solution. LoRA decomposes the locally deployed model into two low-rank matrices, which can then be adjusted in matrices’ ranks based on the local data or device performance demands. This approach decreases both the model training cost and performance requirements and allows each client to customize the model size to its computational capacity and data characteristics. As a result, LoRA makes training and deployment more efficient.
@@
Figure 2: Comparison between traditional FL and LoRA-enabled FL. The Left figure shows the standard FL where clients have full rank $m\times n$ weight matrices. In contrast, the right figure illustrates the LoRA-enabled FL, where clients send low-rank matrices (trained models) in different ranks to reduce communication and computation costs.
@@
Figure. 2 illustrates the differences between traditional centralized training, FedAvg, and FL with LoRA. As shown in the figure, unlike traditional federated learning, FL with LoRA allows each client to customize the rank number based on its own conditions [2]. This flexibility allows the system to adapt to the diverse needs and capabilities of different clients. However, this customization introduces a challenge during the server aggregation process: matrices of different ranks cannot be directly aggregated using traditional methods. To address this, all matrices must be expanded to a uniform dimension for computation.
@@
The current matrix expansion method primarily relies on zero-padding [2], a technique that involves padding smaller matrices with zero elements to match the dimensions of the largest matrix. While this method is straightforward to implement, it introduces significant issues. First, the zero-padded matrices contain a large number of zero elements, leading to unnecessary computational overhead during the aggregation process. Second, padding with zeros can lead to a decline in model performance. These padded zeros do not provide useful information during aggregation. Additionally, the dimensional expansion of matrices can result in a substantial increase in memory usage, which is particularly problematic for resource-constrained devices.
@@
To address the issues posed by heterogeneous matrix ranks in FL, we propose Rank-Based Layer Aggregation (RBLA), a method designed to effectively aggregate matrices of varying ranks. RBLA aims to preserve the high-dimensional features generated during the training process, mitigate the negative impacts of matrix aggregation, and enhance the global model’s convergence speed. The main contributions of our work are as follows:
@@
    We present a toy example demonstrating weight aggregation in LoRA-enabled FL and explain the effect of zero-padding on higher-dimensional features.
@@
    We propose RBLA, a model aggregation method designed to aggregate LoRA models in different ranks.
@@
    Using TensorFlow library, we provide Python implementation of RBLA to compare its performance against state-of-the-art methods across multiple datasets.
@@ 2 Related Work
FL is a decentralized model training method deployable on any device with computational resources and internet connectivity. In practical FL scenarios, clients often possess distinctly different local datasets [17], which may not accurately represent the overall data distribution. Additionally, client device performance is usually heterogeneous [5] due to variations in computational power, memory, and network bandwidth [21]. This heterogeneity can lead to challenges in effectively training and aggregating global model [8, 23], as some clients may contribute less due to their limited resources [19]. The non-IID nature of data can result in a global model that may not generalize effectively across all clients [32].
@@
Various methods have been proposed to address the data non-IID problem. Collins et al. [3] proposed FedRep, which learns a shared representation across clients while allowing each to train personalized local models. Similarly, Ghosh et al. [6] propose the IFCA clustering clients with similar data distributions and optimizing models within each cluster to reduce the impact of data heterogeneity. In the context of personalization, Arivazhagan et al.[1] introduced FedPer, which splits models into shared and personalized layers. FedAvg is used to train shared layers and local-update-personalized layers to obtain better end-side performance. Furthermore, to tackle client drift caused by data heterogeneity, Karimireddy et al. [11] propose SCAFFOLD, an algorithm that uses control variates to correct client drift and perform quicker convergence. To address the issue of device heterogeneity, Khodadadian et al. [13] tackle client’s communication costs by utilizing TD-learning and Q-learning to reduce communication cost overhead, Kumar et al. [16] applied LoRaWAN technology on FL to improve communication efficiency and robustness. Li et al. [19] proposed FedProx, designed to tolerate not-completely-trained local models and use a proximal term to reduce the impact of over-fitted local models. In practical FlaaS involved with multiple IoT devices, the expected local model architecture may differ among client devices due to heterogeneous hardware configurations, tasks, and personal demands [10], [18], [28]. To perform an effective aggregation on a central server with models in different architectures. Wang et al. [26] proposed MaxCommon to aggregate models with different architectures, facilitating collaborative training across various models. To further enhance the flexibility and efficiency of FL in heterogeneous environments, LoRA [9] has emerged as a promising technique. LoRA allows flexible adaption of trainable parameters based on data quality, device performance, etc. Based on LoRA and FL, Yi et al. [29] proposed pFedLoRA, incorporating LoRA in FL to increase model fine-tuning efficiency. Choi et al. proposed HetLoRA [2], dynamically adjusting the client’s rank based on the local model’s quality of fit to local data. Based on LoRA and FL, DFLNet [31] is proposed to improve secure authentication and model convergence in LoRA-enabled networks; LLDPC [27] enhances data transmission reliability in LoRA networks; FDLoRA [25] balances personalized and global learning for large language models (LLMs) while reducing costs; and DP-LoRA [20] ensures differential privacy in FL for LLMs with minimal communication overhead.
@@ 3 Problem Statement
In FL, models must have the same shape to be aggregated on the central server, and zero-padding is one of the most commonly used methods to match dimensions. However, we found that zero-padding introduces structural sparsity, which can slow down convergence. This section presents the impact of zero-padding of trained model weight on the aggregated global model. We specifically analyze how zero-padding affects the sparsity and effectiveness of weighted averaging in the aggregation of LoRA weights [2]. Consider two weight matrices $A_{p\times q}$, $B_{m\times n}$ ($pq<mn$) with corresponding aggregation weights $w_{1}$ and $w_{2}$. To aggregate models with different dimensions, matrix $A$ is padded with zero to match the dimensions of $B$, resulting in the padded matrix $A^{\prime}$ in the shape of $m\times n$. The aggregated model $C_{m\times n}$ is then computed as:
@@
Here, the term ’zero-padding’ introduces non-informative values that negatively impact the overall feature representation, impairing the neuron network’s ability to effectively generalize target features [24]. We highlight the significance of structural sparsity and their impact on neural network capacity and learning dynamics ([24], [30], [7], [12]), which underscore the need to avoid bring structural sparsity during deep feature extraction when aggregating LoRA weights across different dimensions. For example, during the LoRA training process, model $A$ may fail to capture the depth features learned by model $B$ due to insufficient neurons. During aggregation, zero-padding causes the layers of matrix $A$ padded with zeros to dilute the depth features trained by matrix $B$, as these padded layers cannot provide the necessary non-linear transformations [24]. Consequently, zero-padding introduces invalid zero-value information, which dilutes the important depth features learned by model $B$ during the averaging process. This scenario can be illustrated as follows:
@@
Here, $\mathbf{a}_{i,r}$ represents the $r$-th row vector of the zero-padded matrix $A_{i}^{\prime}$ and $\delta_{i,r}$ is an indicator function that equals 1 if the $r$-th row of $A_{i}$ exists and 0 otherwise. This indicates that the more zero-padded is applied to a layer, the more original features of that layer become diluted. Consequently, zero-padding introduces significant issues as it incorporates a substantial amount of zero values, which dilutes effective feature information during aggregation. This dilution reduces the impact of high-dimensional features learned by models with higher dimensions. These zero values, acting as invalid information in the computation, degrade the quality of the aggregated results and lead to a model populated with numerous irrelevant features. In scenarios with skewed data distribution, such as long-tailed distributions where low-rank model clients are assigned fewer classes or data, high-dimensional features learned by deeper neural networks are further diluted by shallower networks during zero-padding weighted average aggregation. This results in additional performance degradation, as zero-padding often diminishes the features from client models with smaller datasets. This issue prevents the aggregated model from fully leveraging all clients’ data and features. Moreover, the structural sparsity introduced by zero-padding imposes invalid linear transformation, which limits the model’s capacity to represent complex patterns and reduces its effectiveness in learning high-dimensional spaces. Additionally, structural sparsity limits the model’s generalization capability and lowers the overall performance. These factors collectively lead to a decline in accuracy and robustness, significantly impacting the models’s effectiveness in practical applications.
@@ 4 RBLA
RBLA is designed to aggregate heterogeneous model weight matrices, bias matrices, and low-rank matrices of different ranks from multiple clients by re-weighting the aggregation weights. Considering matrix sparsity, RBLA first pads all matrices to match the dimensions of the largest LoRA matrix, filling missing entries with a neutral value. It then performs a weighted aggregation for the existing common layer values, either row-wise or column-wise, depending on the presence of none-values and their corresponding aggregation weights. During this procedure, RBLA calculates the aggregation weights based on the common layers, performs a weighted average for shared layers, and preserves the original value of unique stand-alone layers. Figures.4 and 3 illustrate the RBLA aggregation process for models with heterogeneous columns and rows, respectively.
Figure 3: Horizontal aggregation process for heterogeneous LoRA models. Similar to vertical aggregation, RBLA preserves shared layer weights and keeps unique layers.
Figure 4: Vertical aggregation process for heterogeneous LoRA models. Layers of the same color represent common layers. RBLA aggregation preserves the original value of the unique layer and calculates weighted average for common layers.
Algorithm 1 Server Aggregation of RBLA. The server aggregates weights from all clients, accurately handling shared and unique layers through indicator function and weighted aggregation.
@@
The aggregated weight for the $r$-th layer, $\mathbf{C}_{r}$, is computed as:
In this process:
  • $w_{i}$ is the weight coefficient for model $i$.
Applying this method, we can aggregate weights of shared layers while preserving the unique layers of each model to avoid unnecessary structural sparsity.

### v2

@@ Abstract
Federated Learning (FL) is a promising privacy-aware distributed learning framework that can be deployed on various devices, such as mobile phones, desktops, and devices equipped with CPUs or GPUs. In the context of server-based Federated Learning as a Service (FLaaS), FL enables a central server to coordinate the training process across multiple devices without direct access to local data, thereby enhancing privacy and data security. Low-Rank Adaptation (LoRA) is a method that efficiently fine-tunes models by focusing on a low-dimensional subspace of the model’s parameters. This approach significantly reduces computational and memory costs compared to fine-tuning all parameters from scratch. When integrated with FL, particularly in a FLaaS environment, LoRA allows for flexible and efficient deployment across diverse hardware with varying computational capabilities by adjusting the local model’s rank. However, in LoRA-enabled FL, different clients may train models with varying ranks, which poses challenges for model aggregation on the server. Current methods for aggregating models of different ranks involve padding weights to a uniform shape, which can degrade the global model’s performance. To address this issue, we propose Rank-Based LoRA Aggregation (RBLA), a novel model aggregation method designed for heterogeneous LoRA structures. RBLA preserves key features across models with different ranks. This paper analyzes the issues with current padding methods used to reshape models for aggregation in a FLaaS environment. Then, we introduce RBLA, a rank-based aggregation method that maintains both low-rank and high-rank features. Finally, we demonstrate the effectiveness of RBLA through comparative experiments with state-of-the-art methods.
@@ 1 Introduction
Neural Networks (NN) have become a widely applied approach in contemporary Computer Vision (CV) and Natural Language Processing (NLP). Traditionally, model training involves collecting task-specific data and conducting centralized training in high-performance data centers. However, this centralized method raises significant privacy concerns, especially in applications dealing with sensitive data such as health information. Collecting and uploading user data to centralized servers can inherently lead to privacy breaches. To address these privacy challenges, McMahan et al. proposed Federated Learning (FL) in 2017 [24]. FL is a privacy-preserving distributed machine learning framework designed for decentralized neural network training [25]. This algorithm trains models locally on client devices (without sharing raw data), and model updates are then uploaded to a central server for aggregation. This concept has further evolved into Federated Learning as a Service (FLaaS) [17, 7], shown in Figure .1(a), where federated learning capabilities are provided as a cloud-based service [2]. FLaaS simplifies the deployment process by managing the underlying infrastructure, allowing the implementation of federated learning across a diverse range of devices, including smartphones, IoT devices, and edge servers. This flexibility enables organizations to leverage privacy-preserving machine learning on a large scale without the complexities of building and maintaining their own FL systems. Despite these advantages, decentralized local model training in FL typically requires robust client device performance and communication capabilities. In practical FlaaS scenarios, significant performance variations among client devices pose challenges in deploying models with identical structures across all clients.
@@
Recent advancements in computing power and algorithms have led to numerous applications that use large models to process vast amounts of data on mobile devices, providing responses or making decisions. Examples include ChatGPT and Tesla’s autonomous driving technology. However, the inference and training processes of large models are complex and resource-intensive. Additionally, the heterogeneous nature of client devices in FL scenarios, with significant differences in performance, complicates the deployment and fine-tuning of federated large models. To address these issues, Low-Rank Adaptation (LoRA) [11] has emerged as a feasible solution. Figure .1(b) shows LoRA decomposes the locally deployed model into two low-rank matrices, which can then be adjusted in matrices’ ranks based on the local data or device performance demands. This approach decreases both the model training cost and performance requirements and allows each client to customize the model size to its computational capacity and data characteristics. As a result, LoRA makes training and deployment more efficient.
@@
Figure 2: Comparison between traditional FL and LoRA-enabled FL. The left figure shows the standard FL, where clients have full-rank $m\times n$ weight matrices. In contrast, the right figure illustrates LoRA-enabled FL, where clients send low-rank matrices (trained models) with varying ranks to reduce communication and computation costs.
@@
Figure. 2 illustrates the differences between traditional centralized training, FedAvg, and FL with LoRA. As shown in the figure, unlike traditional federated learning, FL with LoRA allows each client to customize the rank number based on its own conditions [3]. This flexibility enables the system to adapt to the diverse needs and capabilities of different clients. However, this customization introduces a challenge during the server aggregation process: matrices of different ranks cannot be directly aggregated using traditional methods. To address this, all matrices must be expanded to a uniform dimension for computation.
@@
The current matrix expansion method primarily relies on zero-padding [3], a technique that involves padding smaller matrices with zero elements to match the dimensions of the largest matrix. While this method is straightforward to implement, it introduces significant issues. First, zero-padded matrices contain a large number of zero elements, leading to unnecessary computational overhead during the aggregation process. Second, padding with zeros can degrade model performance, as the padded zeros do not provide useful information during aggregation. Additionally, the dimensional expansion of matrices can result in a substantial increase in memory usage, which is particularly problematic for resource-constrained devices.
@@
We propose Rank-Based Layer Aggregation (RBLA), a method designed to effectively aggregate matrices of varying ranks to address the issues posed by heterogeneous matrix ranks in FL. RBLA aims to preserve the high-dimensional features generated during the training process, mitigate the negative impacts of matrix aggregation, and enhance the global model’s convergence speed. The main contributions of our work are as follows:
@@
    We present a toy example demonstrating weight aggregation in LoRA-enabled FL and explain the effects of zero-padding on higher-dimensional features.
@@
    We propose RBLA, a model aggregation method specifically designed to aggregate LoRA models of different ranks.
@@
    We provide a Python implementation of RBLA using the TensorFlow library and compare its performance against state-of-the-art methods across multiple datasets.
@@ 2 Related Work
FL is a decentralized model training method deployable on any device with computational resources and internet connectivity. In practical FL scenarios, clients often possess distinctly different local datasets [19], which may not accurately represent the overall data distribution. Additionally, client device performance is usually heterogeneous [6] due to variations in computational power, memory, and network bandwidth [23]. This heterogeneity can lead to challenges in effectively training and aggregating the global model [10, 26], as some clients may contribute less due to their limited resources [21]. The non-IID nature of data can result in a global model that may not generalize effectively across all clients [35].
@@
Various methods have been proposed to address the data non-IID problem. Collins et al. [4] proposed FedRep, which learns a shared representation across clients while allowing each to train personalized local models. Similarly, Ghosh et al. [8] proposed the IFCA, which clusters clients with similar data distributions and optimizes models within each cluster to reduce the impact of data heterogeneity. In the context of personalization, Arivazhagan et al.[1] introduced FedPer, which splits models into shared and personalized layers. FedAvg is used to train shared layers and local-update-personalized layers to obtain better end-side performance. Furthermore, to tackle client drift caused by data heterogeneity, Karimireddy et al. [13] proposed SCAFFOLD, an algorithm that uses control variates to correct client drift and achieve quicker convergence. To address the issue of device heterogeneity, Khodadadian et al. [15] tackled client communication costs by utilizing TD-learning and Q-learning to reduce communication overhead. Kumar et al. [18] applied LoRaWAN technology to FL to improve communication efficiency and robustness. Li et al. [21] proposed FedProx, designed to tolerate not-completely-trained local models and use a proximal term to reduce the impact of over-fitted local models. In practical FlaaS involving multiple IoT devices, the expected local model architecture may differ among client devices due to heterogeneous hardware configurations, tasks, and personal demands [12], [20], and [31]. To perform effective aggregation on a central server with models of different architectures, Wang et al. [29] proposed MaxCommon which facilitates collaborative training across various models. To further enhance the flexibility and efficiency of FL in heterogeneous environments, LoRA [11] has emerged as a promising technique. LoRA allows flexible adaptation of trainable parameters based on data quality, device performance, and other factors. Based on LoRA and FL, Yi et al. [32] proposed pFedLoRA, which incorporates LoRA in FL to increase model fine-tuning efficiency. Choi et al. proposed HetLoRA [3], which dynamically adjusts the client’s rank based on the local model’s quality of fit to the local data. Other approaches, such as DFLNet [34], aim to propose to improve secure authentication and model convergence in LoRA-enabled networks. LLDPC [30] enhances data transmission reliability in LoRA networks. FDLoRA [28] balances personalized and global learning for large language models (LLMs) while reducing costs, and DP-LoRA [22] ensures differential privacy in FL for LLMs with minimal communication overhead.
@@ 3 Problem Statement
In FL, models must have the same shape to be aggregated on the central server, and zero-padding is one of the most commonly used methods to match dimensions. However, we found that zero-padding introduces structural sparsity, which can slow down convergence. This section presents the impact of zero-padding of trained model weights on the aggregated global model. We specifically analyze how zero-padding affects the sparsity and effectiveness of weighted averaging in the aggregation of LoRA weights [3]. Consider two weight matrices $A_{p\times q}$, $B_{m\times n}$ ($pq<mn$) with corresponding aggregation weights $w_{1}$ and $w_{2}$. To aggregate models with different dimensions, matrix $A$ is padded with zeros to match the dimensions of $B$, resulting in the padded matrix $A^{\prime}$ with a shape of $m\times n$. The aggregated model $C_{m\times n}$ is then computed as:
@@
Here, the term ’zero-padding’ introduces non-informative values that negatively impact the overall feature representation, impairing the neuron network’s ability to effectively generalize target features [27]. We highlight the significance of structural sparsity and its impact on neural network capacity and learning dynamics ([27], [33], [9], and [14]), which underscores the need to avoid bringing structural sparsity during deep feature extraction when aggregating LoRA weights across different dimensions. For example, during the LoRA training process, model $A$ may fail to capture the depth features learned by model $B$ due to insufficient neurons. During aggregation, zero-padding causes the layers of matrix $A$ that are padded with zeros to dilute the depth features trained by matrix $B$, as these padded layers cannot provide the necessary non-linear transformations [27]. Consequently, zero-padding introduces invalid zero-value information, which dilutes the important depth features learned by model $B$ during the averaging process. This scenario can be illustrated as follows:
@@
Here, $\mathbf{a}_{i,r}$ represents the $r$-th row vector of the zero-padded matrix $A_{i}^{\prime}$, and $\delta_{i,r}$ is an indicator function that equals 1 if the $r$-th row of $A_{i}$ exists and 0 otherwise. This indicates that the more zero-padded is applied to a layer, the more the original features of that layer become diluted. Consequently, zero-padding introduces significant issues as it incorporates a substantial number of zero values, which dilute effective feature information during aggregation. This dilution reduces the impact of high-dimensional features learned by models with higher dimensions. These zero values, acting as invalid information in the computation, degrade the quality of the aggregated results and lead to a model populated with numerous irrelevant features. In scenarios with skewed data distributions, such as long-tailed distributions where low-rank model clients are assigned fewer classes or data, high-dimensional features learned by deeper neural networks are further diluted by shallower networks during zero-padding weighted average aggregation. This results in additional performance degradation, as zero-padding often diminishes the features from client models with smaller datasets. This issue prevents the aggregated model from fully leveraging all clients’ data and features. Moreover, the structural sparsity introduced by zero-padding imposes invalid linear transformations, which limit the model’s capacity to represent complex patterns and reduce its effectiveness in learning high-dimensional spaces. Additionally, structural sparsity restricts the model’s generalization capability and lowers the overall performance. These factors collectively lead to a decline in accuracy and robustness, significantly impacting the model’s effectiveness in practical applications.
@@ 4 RBLA
Algorithm 1 Server Aggregation of RBLA. The server aggregates weights from all clients by accurately handling shared and unique layers using an indicator function and weighted aggregation.
@@
Figure 3: Horizontal aggregation process for heterogeneous LoRA models: Similar to vertical aggregation, RBLA preserves the weights of shared layers and retains unique layers during the horizontal aggregation process.
RBLA is designed to aggregate heterogeneous model weight matrices, bias matrices, and low-rank matrices of different ranks from multiple clients by re-weighting the aggregation weights. Algorithm. 1 and Algorithm. 2 show the procedures executed on the server and client, respectively. Considering matrix sparsity, RBLA first pads all matrices to match the dimensions of the largest LoRA matrix, filling missing entries with a neutral value. It then performs a weighted aggregation for the existing common layer values, either row-wise or column-wise, depending on the presence of none-values and their corresponding aggregation weights. During this procedure, RBLA calculates the aggregation weights based on the common layers, performs a weighted average for shared layers, and preserves the original value of unique stand-alone layers. Figures 4 and 3 illustrate the RBLA aggregation process for models with heterogeneous columns and rows, respectively.
Figure 4: Vertical aggregation process for heterogeneous LoRA models. Layers of the same color represent common layers. RBLA aggregation preserves the original value of unique layers and calculates a weighted average for common layers.
@@
The aggregated weight for the $r$-th layer, $\mathbf{C}_{r}$, is computed as:
In this process:
  • $w_{i}$ is the weight coefficient for model $i$.
Applying this method, we can aggregate the weights of shared layers while preserving the unique layers of each model to avoid unnecessary structural sparsity.